OpenStudy (anonymous):

Hi can somebody please help me with 3 questions?

OpenStudy (anonymous):

What questions?

OpenStudy (anonymous):

they are all word problems

OpenStudy (anonymous):

Brian is creating a collage on a piece of cardboard that has an area of 110r3 square centimeters. The collage is covered entirely by pieces of paper that do not overlap. Each piece has an area of the square root of r to the fifth power square centimeters. Use the given information to determine an expression for the total number of pieces of paper used.

OpenStudy (anonymous):

Since each piece of paper has the same area ( √r^5 cm^2 ) You need to put √r^5 into exponential notation --> √r^5 = r^5/2 This can be thought of in two ways: a) the square root of r taken to the fifth power or b) finding the square root of the fifth power or r. they are equivalent Multiply it by a variable and set it equal to 100r^3 cm^2. N(r^(5/2) ) =100 r^3 N = 100 r63 / (r^(5/2)) = 100 r^(3 - (5/2)) = 100r^(1/2) N = 100 r^(1/2) = 100√r

OpenStudy (anonymous):

oh okay thank you. heres the other question

OpenStudy (anonymous):

No problem, and okay.

OpenStudy (anonymous):

Julian deposits money into two different savings accounts. He deposits $600 more into Account B than Account A. After a certain amount of time, Account A has earned $180 at a rate of 4%, and Account B has earned $375 at a rate of 5%. How much did Julian initially deposit into each account? A) What equation is used to solve this problem? What does each variable represent? B) Identify the given information. What do you know about P? C) Enter the given information into the equation. D) Solve the equation. E) State the solution.

OpenStudy (anonymous):

375x = 22.5x + 135000 375x - 22.5x = 22.5x + 135000 - 22.5x 352.5x = 135000 352.5x/352.5 = 135000/352.5 x = 382.98

OpenStudy (anonymous):

ok thank you. heres one more

OpenStudy (anonymous):

Create one word problem of your own: rate problem, an interest problem, or a mixture problem. The type of problem you choose must be different from the type you chose to solve in part 1. For example, if you solved the interest problem in part 1 you must create either a rate or mixture problem for part 2. Your word problem must include the following: • word problem (2 pts) • work to solve the word problem (2 pts) • solution (1 pt)

OpenStudy (anonymous):

Wait, There is more to the question you just asked me. he deposited about $382.98 in account A and $982.98 in account B so now you have to add $600 to the amount in account A to get the amount in account B

OpenStudy (anonymous):

oh okay thank you

OpenStudy (amistre64):

are you just posting these to get someone to do the work for you?

OpenStudy (anonymous):

no i just wanted to know how to set it up

OpenStudy (anonymous):

and then i would work it out

OpenStudy (amistre64):

we have to be vigilant here since alot of people try to abuse the nature of this site to just get answers with is all

OpenStudy (anonymous):

yes i understand. im not like that at all. i just need help figuring how to set up the word problem because word problems are my weak point in math

OpenStudy (amistre64):

good luck ;)

OpenStudy (anonymous):

thank you:)